### v1

# Written by AAAI Press Staff1

Abstract
Graph partitioning aims to divide a graph into $k$ disjoint subsets while optimizing a specific partitioning objective. The majority of formulations related to graph partitioning exhibit NP-hardness due to their combinatorial nature. As a result, conventional approximation algorithms rely on heuristic methods, sometimes with approximation guarantees and sometimes without. Unfortunately, traditional approaches are tailored for specific partitioning objectives and do not generalize well across other known partitioning objectives from the literature. To overcome this limitation, and learn heuristics from the data directly, neural approaches have emerged, demonstrating promising outcomes. In this study, we extend this line of work through a novel framework, NeuroCUT. NeuroCUT introduces two key innovations over prevailing methodologies. First, it is inductive to both graph topology and the partition count, which is provided at query time. Second, by leveraging a reinforcement learning based framework over node representations derived from a graph neural network, NeuroCUT can accommodate any optimization objective, even those encompassing non-differentiable functions. Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits resilience to topological modifications.
Abstract
1 Introduction and Related Work
Graph partitioning is a fundamental problem in network analysis with numerous real-world applications in various domains such as system design in online social networks (pujol2009divide), dynamic ride-sharing in transportation systems (tafreshian2020trip), VLSI design (kahng2011vlsi), and preventing cascading failure in power grids (li2005strategic). The goal of graph partitioning is to divide a given graph into disjoint subsets where nodes within each subset exhibit strong internal connections while having limited connections with nodes in other subsets. Generally speaking, the aim is to find somewhat balanced partitions while minimizing the number of edges across partitions.
Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (karypis1997multilevel; karypis1998multilevelk; karypis1999multilevel; andersen2006local; chung2007four). The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel).
In recent years there has been a growing interest in using machine learning techniques to design heuristics for NP-Hard combinatorial problems on graphs. This line of work is motivated from the observation that traditional algorithmic approaches are limited to using the same heuristic irrespective of the underlying data distribution. Neural approaches, instead, focus on learning the heuristic as a function of the data without deriving explicit algorithms (duong2023deep; jung2022learning; nazi2019gap). Significant advances have been made in learning to solve a variety of combinatorial optimization problems such as Set Cover, Travelling Salesman Problem, Influence Maximization etc. (khalil2017learning; gcomb; kool2018attention; joshi2019efficient). The neural architectural designs in these methods are generally suitable for a specific set of CO problems (mazyavkina2021reinforcement) of similar nature. For instance, AM  (kool2018attention) and NeuroLKH (xin2021neurolkh) are targeted towards routing problems. GCOMB (gcomb) and LeNSE (ireland2022lense) aim to solve problems that are submodular in nature. Owing to these assumptions, the architectural designs in these works are not suitable for our problem of $k$-way inductive graph partitioning with versatile partitioning objectives.
The most relevant method to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap). It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation. Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (duong2023deep). Another method solves the multicut problem where the number of partitions is not an input to the problem (jung2022learning). The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function. Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two.
Nevertheless, existing neural approaches that tackle graph partitioning problem suffer three notable drawbacks, which we address.
  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable. As we will illustrate in § 2, the assumption does not always hold in the real-world.
  • Inductivity to partition count: The number of partitions required to segment a graph is an input parameter. Hence, it is important for a learned model to generalize to any partition count without retraining. Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained. Additionally, it’s worth noting that the optimal number of partitions is often unknown beforehand. Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective. Typically, the optimal partition count corresponds to the point where the second-order derivative of the objective function approaches saturation. Therefore, inductivity to the number of partitions becomes an essential requirement for conducting such experiments in a scalable manner.
  • Generalizability to objective functions: Multiple objective functions have been studied in the partitioning literature. The optimal objective function hinges upon the subsequent application in question. Consequently, it holds significance for an algorithmic framework to possess the capacity to encompass a diverse array of objective functions. Unfortunately, prevailing neural frameworks are founded upon presumptions that yield efficacy exclusively within the confines of their designated objective function.
1.1 Contributions:
In this paper, we circumvent the above-mentioned limitations through the following key contributions.
  • Versatile objectives: We develop an auto-regressive, graph reinforcement learning framework, called NeuroCUT, to solve the graph partitioning problem. Diverging from conventional algorithms, NeuroCUT can handle different partitioning objectives. Moreover, unlike other neural methods, NeuroCUT can accommodate diverse partitioning objectives, without the necessity for differentiability.
  • Inductivity to topology and number of partitions: The parameter space of NeuroCUT is independent of the graph size or the partition count. This innovative architecture endows NeuroCUT with the ability to generalize effectively to previously unseen graphs of varying scales and adapt to arbitrary partition count specified during inference.
  • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives. Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions, and demonstrating its capacity to generalize effectively to partition sizes and topological configurations that it has not encountered during training.
2 Problem Formulation
In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Tab. 1 in appendix.
Definition 1 (Graph)
We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{G}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
Definition 2 (Cut)
A cut $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is a partition of $\mathcal{V}$ into two subsets $\mathcal{S}$ and $\mathcal{T}$. The cut-set of $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is the set $(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{T}$ of edges that have one endpoint in $\mathcal{S}$ and the other endpoint in $\mathcal{T}$.
Definition 3 (Graph Partitioning)
Given graph $\mathcal{G}$, we aim to partition $\mathcal{G}$ into $k$ disjoint sets $\mathcal{P}=\{P_{1},P_{2},\cdots,P_{k}\}$ such that the union of the nodes in those sets is equal to $\mathcal{V}$ i.e $\cup_{i=1}^{k}P_{i}=\mathcal{V}$ and each node belongs to exactly one partition.
Partitioning Objective: We aim to minimize/maximize a partitioning objective of the form ${Obj}(\mathcal{G},\mathcal{P})$. A wide variety of objectives for graph partitioning have been proposed in the literature. Without loss of generality, we consider the following four objectives. These objectives are chosen due to being well studied in the literature, while also being diverse from each other.  [^1] Our framework is not restricted to these objectives.
  1. 1.
    $k$-MinCut (saran1995finding): Partition a graph into $k$ partitions such that the total number of edges across partitions is minimized.
[…]
    Here $P_{l}$ refers to the set of elements in $l^{th}$ partition of $\mathcal{P}$ as described in Def.  3 and $\overline{P_{l}}$ refers to set of elements not in $P_{l}$.
  2. 2.
    Normalized Mincut (shi2000normalized): The $k$-minimum cut criteria favors cutting small sets of isolated nodes in the graph. To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
    $\operatorname{Ncut}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}$ (2)
    Here, $\operatorname{vol}\left(P_{l},\mathcal{V}\right):=\sum_{v_{i}\in P_{l},v_{j}\in\mathcal{V}}e(v_{i},v_{j})$.
  3. 3.
    Balanced Cuts (nazi2019gap): Balanced cut favours partitions of equal sizes so an extra term that indicates the squared distance from equal sized partition is added to normalised cuts.
    $\operatorname{Balanced-Cuts}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}+\frac{(|P_{l}|-|\mathcal{V}|/k)^{2}}{|\mathcal{V}|^{2}}$ (3)
    Here, $\operatorname{vol}\left(P_{l},\mathcal{V}\right):=\sum_{v_{i}\in P_{l},v_{j}\in\mathcal{V}}e(v_{i},v_{j})$
  4. 4.
    (Weighted) Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition. We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions. The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
    $\phi(S,\bar{S})$ $=\frac{\operatorname{cut}(S,\bar{S})}{\min(|S|,|\bar{S}|)}$ (4)
    $\operatorname{Sparsest-Cuts}(\mathcal{\mathcal{P}})$ $=\sum_{l=1}^{|\mathcal{P}|}\phi(P_{l},\overline{P_{l}})$ (5)
Problem 1 (Learning to Partition Graph )
Given a graph $\mathcal{G}$ and the number of partitions $k$, the goal is to find a partitioning $\mathcal{P}$ of the graph $\mathcal{G}$ that optimizes a target objective function ${Obj}(\mathcal{G},\mathcal{P})$. Towards this end, we aim to learn a policy $\pi$ that assigns each node $v\in\mathcal{V}$ to a partition in $\mathcal{P}$.
In addition to our primary goal of finding a partitioning that optimizes a certain objective function, we also desire policy $\pi$ to have the following properties:
  1. 1.
    Inductive: Policy $\pi$ is inductive if the parameters of the policy are independent of both the size of the graph and the number of partitions $k$. If the policy is not inductive then it will be unable to infer on unseen size graphs/ number of partitions.
  2. 2.
    Learning Versatile Objectives: To optimize the parameters of the policy, a target objective function is required. The optimization objective may not be differentiable and it might not be always possible to obtain a differentiable formulation. Hence, the policy $\pi$ should be capable of learning to optimize for a target objective that may or may not be differentiable.
Figure 1: Architecture of NeuroCUT. First the initial partitioning of the graph is performed based upon node’s initial embeddings. The initial embeddings are refined using GNN to infuse toplogical information from neighborhood. At each step a node is selected and its partitioned is updated. During the training phase the GNN parameters are updated and hence embeddings are re-computed. During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
Symbol Meaning
$\mathcal{G}$
Graph
$\mathcal{V}$
Node set
$e$
Edge $e\in\mathcal{E}$
$\mathcal{E}$
Edge set
$\mathcal{N}_{v}$
Neighboring nodes of node $v$
$Obj(\mathcal{G},\mathcal{P})$
Objective function based upon graph $\mathcal{G}$ and its partitioning $\mathcal{P}$
$k$
Number of partitions
$\mathcal{P}^{t}$
Partitioning at time $t$
${P}^{t}_{i}$
$i^{th}$ partition at time $t$
$\overline{{P}_{i}}$
Set of nodes that are not in the $i^{th}$ partition
$\mathcal{S}^{t}$
State of system at step $t$
$\operatorname{PART}(\mathcal{P}^{t},v)$
Partition of node $v$ at time $t$
$\mathcal{S}^{t}$
State representation of Partitions and Graph at step $t$
$\mathbf{pos}(v)$
Initial feature representation of node $v$
$\alpha$
Number of anchor nodes for lipschitz embedding
$T$
Length of trajectory
$\pi$
Policy function
Table 1: Notations used in the paper
3 NeuroCUT: Proposed Methodology
Fig. 1 describes the framework of NeuroCUT. For a given input graph $\mathcal{G}$, we first construct the initial partitions of nodes using a clustering based approach. Subsequently, a message-passing Gnn embeds the nodes of the graph ensuring inductivity to different graph sizes. Next, the assignment of nodes to partitions proceeds in a two-phased strategy. We first select a node to change its partition and then we choose a suitable partition for the selected node. Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model. This decoupling allows us to query the model to unseen number of partitions. After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized. The reward us learned through reinforcement learning (RL) (sutton2018reinforcement).
The choice of using RL is motivated two observations. Firstly, cut problems on graphs are generally recognized as NP-hard, making it impractical to rely on ground-truth data, which would be computationally infeasible to obtain. Secondly, the cut objective may lack differentiability. Therefore, it becomes essential to adopt a learning paradigm that can be trained even under these non-differentiable constraints. In this context, RL effectively addresses both of these critical requirements.
We now discuss each of the components in detail.
3.1 Learning Policy $\pi$ as Markov Decision Process
Given a graph $\mathcal{G}$, our objective is to find the partitioning $\mathcal{P}$ that maximizes/minimizes the target objective function $Obj(\mathcal{G},\mathcal{P})$. We model the task of iteratively updating the partition for a node as a Markov Decision Process (MDP) defined by the tuple $(S,\mathcal{A},\rho,R,\gamma)$. Here, $S$ is the state space, $\mathcal{A}$ is the set of all possible actions, $\rho:S\times S\times\mathcal{A}\rightarrow[0,1]$ denotes the state transition probability function, $R:S\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function and $\gamma\in(0,1)$ the discounting factor. We next formalize each of these notions in our MDP formulation.
Initialization. Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions. The graph is clustered into $k$ clusters based upon their positional embeddings (discussed below). We use K-means (macqueen1967some) algorithm for this task. The details of the clustering are present in Appendix B.
Positional Embeddings: Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph. Positional encodings provide an idea of the position in space of a given node within the graph. Two nodes that are closer in the graph, should be closer in the embedding space. Towards this, we use Lipschitz Embedding (bourgain1985lipschitz). Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes. We call them anchor nodes. From each anchor node $i$, the walker starts a random walk (bianchini2005inside) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$. Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$. Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
$\vec{r}_{i}=c\tilde{\mathbf{W}}\vec{r}_{i}+(1-c)\vec{e}_{i}$ (6)
Eq. 6 describes the random walk starting at node $i$. In vector $\vec{e}_{i}\in\mathbb{R}^{|V|\times 1}$, only the $i^{th}$ element (the initial anchor node) is $1$, and the rest are set to zero. We set $\mathbf{W}_{ij}=\frac{1}{degree(j)}$ if edge $e_{ij}\in\mathcal{E}$, 00 otherwise. The random walk with restart process is repeated for $\beta$ iterations, where $\beta$ is a hyper-parameter. Here $\vec{e}_{i}\in\mathbb{R}^{|V|\times 1}$ and $c$ is a scalar.
Based upon the obtained random walk vectors for the set of anchor nodes $\mathcal{A}$, we embed all nodes $u\in\mathcal{V}$ in a $\alpha$-dimensional feature space:
$\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$ (7)
State: The state space characterizes the state of the system at time $t$ in terms of the current set of partitions $\mathcal{P}^{t}$. Intuitively the state should contain information to help our model make a decision to select the next node and the partition for the node to be assigned. Let $\mathcal{P}^{t}$ denote the status of partitions at time $t$ wherein a partition $P^{t}_{i}$ is represented by all nodes belonging to $i^{th}$ partition. The state of the system at step $t$ is defined as
$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{pos}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (8)
Here state of each partition is represented by the collection of positional embedding of nodes in that partition.
Action: Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes. The first phase consists of identifying a node to update its partition. Instead of arbitrarily picking a node, we propose to prioritize selecting nodes for which the new assignment is more likely to improve the overall partitioning objective. In comparison to a strategy that arbitrarily selects nodes, the above mechanism promises greater improvement in the objective with less number of iterations. In the second phase, we calculate the score of each partition $\mathcal{P}$ with respect to the selected node from the first phase and then assign it to one of the partitions based upon the partitioning scores. We discuss both these phases in details below.
Phase 1: Node Selection Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$. Our proposed formulation involves selecting a node $v$ at step $t$ belonging to partition $\operatorname{PART}(\mathcal{P}^{t},v)$ and then assigning it to a new partition. The newly assigned partition and the current partition of the node could also be same.
Towards this, we design a heuristic to prioritize selecting nodes which when placed in a new partition are more likely to improve the overall objective value. A node $v$ is highly likely to be moved from its current partition if most of its neighbours are in a different partition than that of node $v$. Towards this, we calculate the score of nodes $v\in\mathcal{V}$ in the graph as the ratio between the maximum number of neighbors in another partition and the number of neighbors in the same partition as $v$. Intuitively, if a partition exists in which the majority of neighboring nodes of a given node $v$ belong, and yet node $v$ is not included in that partition, then there is a high probability that node $v$ should be subjected to perturbation. Specifically the score of node $v$ at step $t$ is defined as:
$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in\mathcal{P}^{t}-\operatorname{PART}(\mathcal{P}^{t},v)}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p|}}}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=\operatorname{PART}(\mathcal{P}^{t},v)|}}$ (9)
Further, a node having a higher degree implies it has several edges associated to it, hence, an incorrect placement of it could contribute to higher partitioning value. Hence, we normalize the scores by the $degree$ of the node.
Phase 2: Inductive Neural Method for Partition Selection Once a node is selected, the next phase involves choosing the new partition for the node. Towards this, we design an approach empowered by Graph Neural Networks (GNNs) which enables the model to be inductive with respect to size of graph. Further, instead of predicting a fixed-size score vector (nazi2019gap) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too. We discuss both above points in section below.
Message passing through graph neural network: To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (gcn; velivckovic2017graph; hamilton2017inductive). GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{pos}(u)$ using eq. 6. We perform $L$ layers of message passing to compute representations of nodes. To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\mathrm{mean}_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (10)
where $\mathbf{h}_{u}^{(l)}$ is the node embedding in layer $l$. $\mathbf{W}^{l}_{1}$ and $\mathbf{W}^{l}_{2}$ are trainable weight matrices at layer $l$.
Following $L$ layers of message passing, the final node representation of node $u$ in the $L^{th}$ layer is denoted by $\mathbf{h}_{u}^{L}\in\mathbb{R}^{d}$. Intuitively $\mathbf{h}_{u}^{L}$ characterizes $u$ using a combination of its own features and features aggregated from its neighborhood.
Scoring partitions: Recall from eq. 8, each partition at time $t$ is represented using the nodes belonging to that partition. Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$. In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap), the proposed design choice makes the model inductive to the number of partitions. Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
Having obtained the transformed node embeddings through a Gnn in Eq. 10, we now compute the (unnormalized) score for node $v$ selected in phase 1 for each partition $p\in\mathcal{P}^{t}$ as follows:
$\operatorname{PartScore}$ $(p,v){=}\operatorname{AGG}(\{\operatorname{MLP}(\sigma(h_{v}|h_{u}))$ (11)
$\quad\quad\quad\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p\})$
The above equation concatenates the selected node $v$’s embedding with its neighbors $u\in\mathcal{N}(v)$ that belong to the partition $p$ under consideration. In general, the strength of a partition assignment to a node is higher if its neighbors also belong to the same partition. The above formulation surfaces this strength in the embedding space. The concatenated representation $(h_{v}|h_{u})\forall u\in\mathcal{N}(v)$ is passed through an MLP that converts the vector into a score (scalar). We then apply an aggregation operator (e.g., mean) over all neighbors of $u$ belonging to $\mathcal{P}^{t}$ to get an unnormalized score for partition $p$. Here $\sigma$ is an activation function.
To compute the normalized score at step $t$ is finally calculated as softmax over the all partitions $p\in\mathcal{P}^{t}$ for the currently selected node $v$. Mathematically, the probability of taking action $a^{t}{=}p$ at time step $t$ at state $S^{t}$ is defined as:
$\pi((a^{t}{=}p)/S^{t})=\frac{exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (12)
During the course of trajectory of length $T$, we sample action $a^{t}\in\mathbb{R}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
State Transition: After action $a^{t}$ is applied at state $S^{t}$, the state is updated to $S^{t+1}$ that involves updating the partition set $\mathcal{P}^{t+1}$. Specifically, if node $v$ belonged to $i^{th}$ partition at time $t$ and its partition has been changed to $j$ in phase 2, then we apply the below operations in order.
$P^{t+1}_{i}\leftarrow P^{t}_{i}\setminus v\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup v$ (13)
Reward: Our aim is to improve the value of the overall partitioning objective. One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$. Here, the denominator term ensures that the model receives a significant reward for even slight enhancements when dealing with a small objective value. However, this definition of reward focuses on short-term improvements instead of long-term. Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term. The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term). Mathematically,
$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}\vspace{-0.1in}$ (14)
where $T$ is the length of the horizon and $\gamma\in(0,1]$ is a discounting factor (hyper-parameter) describing how much we favor immediate rewards over the long-term future rewards.
The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable. This is an advantage over existing neural method (nazi2019gap) where having a differentiable form of the partitioning objective is a pre-requisite.
3.2 Policy Loss Computation and Parameter Update
Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others. Towards this, we use REINFORCE gradient estimator (williams1992simple) to optimize the parameters of our policy network. Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$. Towards this end, we define a reward function $J(\pi_{\theta})$ as:
$J(\pi_{\theta})=\mathbb{E}\big{[}\sum_{t=0}^{T}\left(D^{t}\right)\big{]}$ (15)
We, then, optimize $J(\pi_{\theta})$ as follows:
${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$ (16)
3.3 Training and Inference
For a given set of training graphs, we optimize the parameters of the policy network $\pi_{\theta}$ for $T$ steps. Subsequently, given a new graph and partition objective, we adapt the policy parameters $\pi_{\theta}$ using Eq. 16. Note that the trajectory length $T$ is not kept very large to avoid the long-horizon problem.
During inference, we compute the initial node embeddings, obtain initial partitioning and then run the forward pass of our policy to improve the partitioning objective over time.
3.4 Time complexity
The time complexity of NeuroCUT during inference is
$\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$. Here $\alpha$ is number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Further, for sparse graphs $\mathcal{E}=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$. In comparison to spectral clustering, which is a common graph partitioning algorithm having time complexity of $\mathcal{O}(|\mathcal{V}|^{3})$, the proposed technique’s complexity is much lower. This can also be attributed to the fact that NeuroCUT exploits the sparse nature of graphs unlike spectral clustering that performs matrix operations of $\mathcal{O}(|\mathcal{V}|^{3})$. For detailed derivation please see appendix  C.
[…]
Figure 2: Results of all methods on five datasets through radar plots. The color shade represents identity of each method. The legend in the first subplot is common for all plots. All values (for each partitioning objective) are normalized by dividing by the worst (maximum) cut value obtained on that dataset. The smaller the value, i.e, closer to the center, better is the performance. Further, when two polygons are overlapping, but not fully contained within each other, in general, the smaller polygon tends to indicate a better holistic performance across all objectives. Fig (a) represents cut values averaged across five datasets. Fig (b-f) depict results on each of the five datasets. The neural state-of-the-art GAP method produces the most inferior performance among considered baselines. Further in several cases it generates ”NaN” values (error). Our method NeuroCUT  either outperforms the baselines or produces the second best results while being close to the best one in all cases. Moreover, NeuroCUT  is the most robust and does not produce the worst cut values. The neural methods are trained and tested with the same number partitions ($k=5$). Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
4 Experiments
In this section, we demonstrate the efficacy of NeuroCUT against state-of-the-art methods and establish that:
  • Efficacy and robustness: NeuroCUT consistently produces the best or the second best results in terms of the average performance over all objective functions. This establishes the robustness of NeuroCUT.
  • Inductivity: As one of the major strengths, unlike the state-of-the-art neural model GAP (nazi2019gap), NeuroCUT is inductive on the number of partitions and, consequently, generalizes well over unseen number of partitions.
4.1 Experimental Setup
Datasets:
We use five datasets for our experiments among which one is synthetic. They are as follows:
  • Cora and Citeseer (sen2008collective): These are citation networks. In these, a $k$-way cut could represent the partitioning of publications into $k$ clusters while minimizing the citations across the clusters. Each cluster might correspond to a distinct research field, shedding light on interdisciplinary connections or isolating specialized domains.
  • Facebook (leskovec2012learning): In social network analysis, a $k$-way cut can delineate communities of users with minimal inter-community connections. This partitioning unveils hidden social circles and provides insights into how different groups interact within the larger network. The datasets have multiple connected components, we work with the largest connected component to make sure we can evaluate proper cuts.
  • Harbin (li2019learning): In this road network, a $k$-way cut could entail partitioning road segments into $k$ subsets to optimize traffic flow between them. This partitioning aids in urban planning by identifying key intersections and routes that minimize congestion.
  • Stochastic Block Models (SBM) (abbe2017community): This synthetic network is generated by SBM where $k$-way cuts reveal the underlying community structure. This helps to validate our method by comparing the identified partitions with the ground truth communities.
Baselines:
We compare our proposed method NeuroCUT with the state-of-the-art neural method called GAP (nazi2019gap). It allows balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut. We also use two popular non-neural baselines: hMetis (karypis1997multilevel) and spectral clustering (ng2001spectral).
Other settings:
We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments. We used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$. We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$. For computing discounted reward in RL, we use discount factor $\gamma=0.99$. We set the length of trajectory $T$ during training as $2$. At time step $t$, the rewards are computed from time $t$ to $t+T$ and parameters of the policy $\pi_{\theta}$ are updated. The default number of anchor nodes for computing positional embeddings is set to $35$.
4.2 Results on Transductive Setting
We compare our proposed method NeuroCUT against all the baselines where the neural models are trained and tested with the same number of partitions ($k=5$). Fig. 2 presents the results for all datasets and all methods. The results are depicted using spider plots where the smaller the value i.e closer to the center, better is the performance. Further, in the plots, if two polygons are overlapping, but not fully contained within each other, in general the smaller polygon tends to indicate a better holistic performance across all objectives (heckert2002handbook). NeuroCUT outperforms the baselines or produces the next to optimal results. Moreover, NeuroCUT is the most robust among all the methods and does not produce the worst cut values in any of the settings. Overall, the performance of NeuroCUT is the most consistent among all the methods. Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
NeuroCUT incorporates causality during inference and this leads to the robust performance observed in fig. 2. Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in GAP (nazi2019gap). We observe that the non-neural methods fail to generalize well across all objective functions since they do not learn the base heuristic as a function of the objective and the dataset. Finally, we note that in SBMs the task is easier as there are already $5$ clusters which aligns well with the the desired number of partitions, which is also $5$.
4.3 Results on Inductive Setting
Dataset Method $normalised$ $sparsest$ $balanced$ $kmin$
Cora hMetis 0.772 0.772 3.177 0.097
Spectral 0.723 1.503 1.845 0.008
NeuroCUT 0.591 2.451 0.981 0.022
CiteSeer hMetis 0.415 0.415 1.554 0.056
Spectral 0.328 1.094 0.792 0.003
NeuroCUT 0.363 1.060 0.457 0.031
SBMs hMetis 3.259 3.260 65.955 0.393
Spectral 2.495 2.513 50.805 0.242
NeuroCUT 3.241 5.505 45.602 0.038
Table 2: Inductive results based on the number of partitions of all methods on all six datasets. The best ones are shown by green and the second best ones are in gray. Lower values are better. Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in ten out of twelve settings. NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$). GAP is not inductive in terms of number of partitions and thus the results are not present here.
We show that NeuroCUT generalizes to unknown number of partitions. We compare it against only the non-neural baselines since GAP does not generalize to unseen number of partitions. Here, NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$). Table 2 shows the results. Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in most of the settings. Note that while the non-neural methods need to be re-run for the unseen partitions, NeuroCUT only runs forward pass to produce the results on an unseen partition size.
4.4 Generalization to Topological changes in graphs
In this section, we analyze the performance of NeuroCUT to topological changes in the graph. Specifically, once a model is trained, we aim to evaluate its performance when the topology of the graph changes which is a common scenario in dynamic networks such as social networks where connections are formed/deleted over time.
Towards this first we train NeuroCUT on a given graph. Next, we add/remove $x\%$ percentage of edges in the graph. Then we predict the cuts on the perturbed graph. To predict cuts for the perturbed graph, our model NeuroCUT only needs to just run a forward pass over it.
Tables  3,4,5,6 show the results on two datasets, namely, Cora and Facebook for different partitioning objectives at different degrees of perturbations. In almost all settings, our method NeuroCUT outperforms the baselines. The only neural baseline i.e GAP had to be retrained on perturbed graph and still produced NaN values(indicated by -) in most of the cases. These results show that our proposed method NeuroCUT is more robust in comparison to existing neural methods at different levels of perturbations. Additionally, in several cases it outperforms non-neural methods which were directly optimized on perturbed graphs.
[…]
Cora hMetis No 0.351 0.355 0.399 0.435 0.506
Spectral No 0.346 0.442 0.479 0.424 0.490
GAP Yes 0.512 0.500 0.483 0.808 0.664
NeuroCUT Yes 0.321 0.307 0.364 0.428 0.569
Facebook hMetis No 1.112 1.113 1.126 1.126 1.131
Spectral No 2.034 1.676 1.918 1.92 1.679
GAP Yes - - 0.482 - -
NeuroCUT Yes 0.487 0.459 0.463 0.494 0.864
Table 3: Perturbing $x\%$ Edges - Normalised Cuts
[…]
Cora hMetis No 1.442 1.464 1.649 1.819 2.169
Spectral No 0.795 1.058 1.224 0.957 1.452
GAP Yes 2.029 1.978 1.915 3.167 2.816
NeuroCUT Yes 1.015 1.317 1.479 1.546 2.053
Facebook hMetis No 51.669 52.937 52.110 52.962 52.236
Spectral No 5.625 4.722 5.591 5.591 4.867
GAP Yes - - 11.886 - -
NeuroCUT Yes 3.528 3.516 3.745 4.060 5.298
Table 4: Perturbing $x\%$ Edges - Sparsest Cuts
[…]
Cora hMetis No 0.352 0.355 0.399 0.435 0.506
Spectral No 1.089 1.223 0.800 1.207 0.841
GAP Yes 0.536 0.529 0.515 0.870 0.681
NeuroCUT Yes 0.526 0.544 0.613 0.651 0.836
Facebook hMetis No 1.112 1.113 1.126 1.127 1.131
Spectral No 2.818 2.459 2.697 2.697 2.456
GAP Yes - - 0.506 - -
NeuroCUT Yes 0.831 0.793 0.800 0.855 0.921
Table 5: Perturbing $x\%$ Edges - Balanced Cuts
[…]
Cora hMetis No 0.071 0.072 0.080 0.087 0.101
Spectral No 0.002 0.001 0.054 0.001 0.064
GAP Yes 0.099 0.097 0.094 0.114 0.126
NeuroCUT Yes 0.010 0.013 0.014 0.018 0.021
Facebook hMetis No 0.199 0.203 0.201 0.203 0.201
Spectral No 0.000 0.000 0.000 0.000 0.000
GAP Yes 0.024 0.026 0.044 0.026 0.027
NeuroCUT Yes 0.000 0.000 0.000 0.000 0.000
Table 6: Perturbing $x\%$ Edges - KMin Cuts
[…]
Figure 3: Results on the initial and the final cut values at $k=5$. It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function. Subsequently, there is a significant difference between the initial and final cut values.
4.5 Ablation Studies
Initial vs Final Cut Values. NeuroCUT takes a warm start by partitioning nodes based on clustering over their Lipschitz embeddings, which are subsequently fine-tuned auto-regressively through the reinforcement learning based framework. How much does the partitioning objective improve since the initial clustering? Figure 3 sheds light on this question. Specifically, it shows the difference between initial cut value due after clustering and the final cut value from the partitions produced by our method. We note that there is a significant difference between the initial and final cut values, which essentially shows the effectiveness of NeuroCUT.
[…]
Figure 4: NeuroCUT results with heuristics for node selection in Phase 1 for $k=5$. Our heuristic finds better cuts than a random selection algorithm for two objective functions: Normalized Cut and KMin Cut.
Impact of Node Selection Procedures. Figure 4 shows how our proposed heuristic for node selection helps in finding the near-optimal cuts for any given objective function. We compare and show that a simpler node selection where we select all the nodes one by one in a random order produces substantially inferior results. Furthermore, the difference lies not only in final cut-values but also we require much less perturbations.
In App. D of appendix, we present additional results on variation of parameters of NeuroCUT.
5 Conclusion
In this paper, we have introduced a novel graph reinforcement learning framework NeuroCUT to effectively address the graph partitioning problem. A distinctive feature of our framework is its adaptability to diverse partitioning objectives that may or may not be differentiable, which sets it apart from conventional algorithms. Our approach has also shown the ability to generalize on unseen graphs as well as unseen number of partitions. The efficacy of our approach is empirically validated through an extensive evaluation on six datasets with four graph partitioning objectives. Notably, our method shows remarkable performance when compared to both non-neural and neural state-of-the-art techniques, proving its competence in both inductive and transductive settings. In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Method $normalised$ $sparsest$ $balanced$ $kmin$ AVG DIFF MAX DIFF
ACTDIFF ACTDIFF ACTDIFF ACTDIFF
Cora 2485 5069 hMetis 0.343   0.115 1.412   0.061 0.345   0.000 0.069   0.063 0.060 0.115
Spectral 0.370   0.142 1.351   0.000 1.165   0.820 0.006   0.000 0.240 0.820
GAP 0.505   0.277 2.074   0.723 0.536   0.191 0.095   0.089 0.320 0.723
NeuroCUT 0.228   0.000 1.351   0.000 0.570   0.225 0.011   0.005 0.057 0.225
CiteSeer 2120 3679 hMetis 0.198   0.025 0.673   0.227 0.198   0.000 0.038   0.037 0.072 0.227
Spectral 0.173   0.000 0.446   0.000 0.929   0.731 0.001   0.000 0.182 0.731
GAP 0.267   0.094 0.795   0.349 0.312   0.114 0.043   0.042 0.150 0.349
NeuroCUT 0.211   0.038 0.712   0.266 0.296   0.098 0.036   0.035 0.109 0.266
SBMs 500 5150 hMetis 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
Spectral 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
GAP - - - 0.035   $0$
NeuroCUT 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
Facebook 1034 26749 hMetis 1.115   0.858 52.319   48.198 1.115   0.143 0.200   0.197 12.349 48.198
Spectral 1.670   1.413 4.722   0.601 2.459   1.487 0.003   0.000 0.875 1.487
GAP - - - 0.025   0.022 - -
NeuroCUT 0.257   0.000 4.121   0.000 0.972   0.000 0.015   0.012 0.003 0.012
Harbin 6235 15205 hMetis 0.058 0.000 0.184   0.000 0.058   0.000 0.011   0.000 0.000 0.000
Spectral 0.104   0.046 0.348   0.164 0.403   0.345 0.015   0.004 0.140 0.345
GAP 0.175   0.117 0.551   0.367 0.269   0.211 0.017   0.006 0.175 0.367
NeuroCUT 0.092   0.034 0.286   0.102 0.099   0.041 0.018   0.007 0.046 0.102
Table 7: Results of all methods on all five datasets. “ACT” denotes the actual values of the measures and “DIFF” shows the difference with the best value for the dataset. “AVG DIFF” and “MAX DIFF” present the average and the maximum of the differences over all the measures respectively. The best average differences are shown by green and the second best ones are in gray. The worst result, i.e., the maximum difference is shown in red. The neural state-of-the-art GAP method produces the most inferior performance among considered baselines. Our method NeuroCUT either outperforms the baselines (green) or produces the second best (gray) results while being close to the best one in all cases. Moreover, NeuroCUT is the most robust and never produces the worst maximum difference (i.e., in red). The neural methods are trained and tested with the same number partitions ($k=5$).
Appendix A Result for all datasets in transductive setting
In Table  7 we present the results obtained by all methods on different datasets. These correspond to radar plot obtained in fig. 2 in main paper .
Appendix B Clustering for Initialization
As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions. Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its positional node embedding i.e $pos(u)\;\forall u\in\mathcal{V}$. Further, we used $L_{\inf}$ norm as the distance metric for clustering.
Appendix C Time complexity of NeuroCUT
We derive the time complexity of forward pass of NeuroCUTas discussed in sec. 3.4 in main paper.
(1) First the positional embeddings for all nodes in the graph are computed. This involves running RWR for $\alpha$ anchor nodes for $\beta$ iterations(Eq.  6). This takes $\mathcal{O}(\alpha\times\beta)$ time. (2)Next, GNN is called to compute embeddings of node. In each layer of GNN a node $v\in\mathcal{V}$ aggregates message from $d$ neighbors where $d$ is the average degree of a node. This takes $\mathcal{O}(\mathcal{V})$ time. This operation is repeat for $L$ layers. Since $L$ is typically 1 or 2 hence we ignore this factor. (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq. 9. This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time. (4) Finally for the selected node, its partition has to be determined using eq. 11 and  12. This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score. Steps 2-4 are repeated for $T^{\prime}$ iterations. Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Since $|\mathcal{V}|\times d\approx|\mathcal{E}|$, hence complexity is $o(|\mathcal{E}|\times|\mathcal{V}|\big{)}$ Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
Appendix D Impact of number of anchor nodes
In this section we analyze the impact of number of anchor nodes chosen for computing positional embeddings of nodes in the graph $pos(u)$. Fig. 5 shows the results where $Y$-axis shows the normalized cut value and $X$-axis shows the number of anchor nodes. The best results are achieved at the number of anchor nodes being 35 except in Citeseer. The other variation in the number of anchor nodes produces similar results often.
Figure 5: Results on varying the number of anchor nodes on the normalized cut value produced by NeuroCUT.